# The vertical structure of debris discs and the role of disc gravity

### v1

@@ 5.2 Disks Less Massive than the Planet ($M_{d}/m_{p}\lesssim 1$)
Stage 1 $[0\leq t\lesssim T_{\dot{\Omega}_{p}}]$: At early times, i.e., before the planet completes one full nodal precession cycle, the disk’s morphological evolution qualitatively resembles that of a massless disk. Indeed, as seen in Figures 9(i) and (ii), a spiral bending wave is launched at $a_{\rm in}$, which propagates outward as a leading pattern while gradually wrapping around the star. This manifests as a double-armed spiral in the top-down images and as a warp in the edge-on views, consistent with “Stage 1” evolution for $M_{d}=0$ disks (Section 5.1). However, this description is valid only within a certain radial range and is subject to minor modifications, as outlined below.
@@
Stage 2 $[t\sim T_{\dot{\Omega}_{p}}]$: By the time the planet has completed one full precession cycle, the warp extends to $a_{\rm res}$, and the disk’s morphology exhibits significant deviations from that of a massless disk; see Figure 9(iii). First, the inner tilted region of the disk no longer transitions smoothly into the outer parts but instead does so rather abruptly. This creates the appearance of a “broken” disk, in which the region interior to the resonance appears separated from the outer parts. This can be understood by noting that planetesimal inclinations are maximized when nodally aligned with the planet in the inner region, but anti-aligned in the outer region (Figure 5(b)). This disconnect is most evident around the resonance, where inclinations are higher, especially compared to the inclinations toward the disk’s outer edge (recall that $I_{d}(a_{d})\propto a_{d}^{-7/2}$ at $a_{d}\gtrsim a_{\rm res}$; Equation 32).
@@
Stage 3 $[t\gtrsim T_{\dot{\Omega}_{p}}]$: Further into the evolution, the disk settles into and maintains a quasi-equilibrium state, remaining largely unaffected by the continued growth of inclinations at the resonance; see panels (iv)–(vi) in Figure 9. Indeed, the disk structure – including the curly “X”-pattern – remains invariant, coprecessing with the planet’s ascending node as a rigid body. As this occurs, the broken, warp-like feature stalls near $a_{d}\approx a_{\rm res}$, persisting as a long-lived feature.
@@ 5.3 Disks More Massive than the Planet ($M_{d}/m_{p}\gtrsim 1$)
Stage 1 $[0\leq t\lesssim T_{\dot{\Omega}_{p}}]$: At early times, the disk departs from its initial razor-thin configuration by developing a warp-like feature near its inner edge at $a_{\rm in}=30$ au (Figure 10(i)). This feature, however, is not directly analogous to the warp that develops in massless disks (Section 5.1). The key difference is that the location of the warp does not propagate outward over time. Instead, the maximum vertical displacement from the initial midplane gradually increases while remaining at an in-plane radius of $\approx a_{\rm in}$, such that $|Z_{\rm max}|\sim\sin[I_{d}(a_{\rm in})]$ (Figures 10(i)–(iii)). This effect is amplified as $I_{d}(a_{\rm in})\rightarrow 90^{\circ}$ in this specific setup (Figure 5(c)), although similar behavior would occur even if $I_{d}(a_{\rm in})$ were smaller. [^12] For instance, for a system with a larger $M_{d}/m_{p}$ and/or a smaller $a_{p}/a_{\rm in}$, such that $A_{d,p}/A_{p}$ at $a_{d}\approx a_{\rm in}$ is relatively larger than unity, rather than comparable to as in Model C; see the green curve in Figure 2. Additionally, this structure precesses as a whole while maintaining its symmetry around the planetary orbit.
@@
Finally, at $a_{d}\gtrsim a_{\rm in}$, the disk remains mostly undisturbed, preserving its flat structure. This is because, although planetesimals may have already attained their maximum inclinations, these decrease steeply with distance and remain small, $I_{m}(a_{d})\propto a_{d}^{-7/2}\ll 1^{\circ}$, with $\Delta\Omega(a_{d})\approx{\rm cte}$; see, e.g., panel (ii) in Figures 5(c) and 10.
Stage 2 $[t\gtrsim T_{\dot{\Omega}_{p}}]$: Once the planet has completed more than one precession cycle, the warp-like feature begins to propagate outward, akin to the scenario with $M_{d}=0$ (Figures 10(iv)–(vi)). This occurs as the planetesimal inclination profile $I_{d}(a_{d})$ develops oscillations along $a_{d}$, and the ascending nodes $\Delta\Omega(a_{d})$ become phase-mixed in the inner disk regions; see also Figure 5(c). This description also explains the appearance of spiral patterns in the disk when viewed top-down, particularly near $a_{\rm in}$, where inclinations are larger.
@@ 5.4.4 Variation of the planetary semimajor axis, $a_{p}$
$\phi_{1}^{c}(a_{p}/a_{\rm in})\frac{M_{d}}{m_{p}}\sqrt{\frac{a_{\rm in}}{a_{p}}}=\mathrm{cte}.$ (45)
@@ 6 Discussion
To address the questions raised in Section 1.3 regarding the extent to which planetary inferences are compromised by the disk mass, we first discuss the implications of our findings for interpreting warped structures (Section 6.1), followed by vertical scale heights and density profiles (Section 6.2). Finally, we discuss the implications of our study for constraining the total masses of debris disks (Section 6.3), before reviewing the key assumptions and limitations of our model (Section 6.4).
@@ 6.1.2 Application to HD 110058
[…]
@@
Figure 11: The mass of a hypothetical planet $m_{p}$ as a function of its semimajor axis $a_{p}$ capable of producing a warp in the HD 110058 debris disk at $a_{\rm warp}=50$ au. The solution with a massless disk is depicted by the blue dashed curve ($M_{d}=0$). For the massive disks, two families of solutions are presented (assuming $p=3/2$; Eq. (1) and Section 6.1.1): the black and gray contours correspond to “Family 1” (i.e., $a_{\rm warp}=a_{\rm res}$) and “Family 2” (i.e., $a_{\rm warp}<a_{\rm res}$), respectively, with different contours representing different values of $M_{d}$ (in $M_{\earth}$; as indicated). Color-shaded regions are excluded: (i) the gray region represents where the disk would be too massive; (ii) the green region corresponds to inclination excitation timescales at the resonance that are longer than the system age; (iii) the yellow region indicates where the planet would have carved an inner disk edge larger than the one observed (Eq. 3); and (iv) the olive region represents where $m_{p}\geq M_{c}/10$. The remaining white (unshaded) area represents the allowed parameter space where all of the above conditions are satisfied. Note that the red region is where the presence of companions is ruled out with $>68\%$ confidence (based on direct imaging and Gaia astrometry; Stasevic et al. (2023)). See the text (Section 6.1.2) for details.
@@
The results of Figure 11 are promising, in the sense that (planetary) companions responsible for the warp in HD 110058 could be more massive than currently thought (especially at $\sim 10$ au separations); therefore, potentially detectable in future observations. Currently, direct imaging and proper motion anomaly have ruled out giant planets outside the HD 110058 disk (namely, with $m_{p}\gtrsim 8M_{J}$ at $\gtrsim 50$ au; Brandt (2021); Wahhaj et al. (2013); Meshkat et al. (2015); Kervella et al. (2022)), and there is a likelihood of $<68\%$ of detecting any companion interior to the disk with a mass smaller than $\sim 50M_{J}$ (Stasevic et al., 2023, see the red-shaded region in Figure 11).
@@ 6.1.3 The case of $\beta$ Pictoris
Figure 13: Vertical characteristics of the debris disks in each of models A, B, and C (Table 1). For each model, results correspond to the final snapshot time shown in Figure 5. Panel (A): Mass-weighted distributions of planetesimal inclinations, computed using Equation (18) and $p=3/2$, shown as histograms with $100$ equal-sized bins. For reference, the black vertical line represents twice the planetary inclination $2I_{p}(0)$. Note that in Model C, nearly $78\%$ of planetesimals have inclinations of $1^{\circ}$ or less. Panel (B): Normalized vertical density profiles as a function of distance from the disk midplane $Z_{\rm mid}$. Results are averaged over a radial range of ±$\pm 2.5$ au centered on the reference projected radius $R_{0}\approx 49.1$ au (equal to $\sqrt{a_{\rm in}a_{\rm res}}$ in Model B), using a resolution of $dZ_{\rm mid}=0.1$ au. The inset presents a zoomed view around the midplane. Regardless of disk mass, the non-Rayleigh nature of the inclination distributions and the non-Gaussian shape of the density profiles are evident. For $M_{d}/m_{p}=0.1$, the inclination distribution is bimodal (apart from the peak at $I_{d}=90^{\circ}$), and the density profile resembles a sum of two Gaussians with different widths, indicative of dynamically cold and hot populations. See the text (Section 6.2.1) for details.
@@ 6.2.2 Aspect ratios: radial dependence and implications
Second, the radial profile of a disk’s aspect ratio – whether constant or decreasing with $R$ – can provide insights into the effects of disk gravity. In principle, this could constrain the parameters of a potential planet as a function of $M_{d}$ by making use of Equations (49)–(51), or conversely, offer a dynamical estimate of the disk mass if a planet is already known. However, by and large for simplicity, most model fits to observations typically assume a Gaussian vertical profile with a distance-independent aspect ratio, $\mathcal{H}(R)={\rm cte}$ (e.g., Kennedy et al., 2018; Daley et al., 2019; Vizgan et al., 2022; Terrill et al., 2023; Matrà et al., 2025). As a result, this approach cannot currently be pursued, and future modelling efforts should aim to account and fit for the radial dependence of vertical scale heights.
We note that a recent study by Han et al. (2025) explored this avenue using a non-parametric fitting procedure. They found that 8 out of 9 considered disks are consistent with a distance-independent aspect ratio or an $\mathcal{H}(R)$ which increases with radius in their millimeter dust continuum. Interestingly, the debris disk around HD 15115 (MacGregor et al., 2019, not known to have companions) stands out due to its decreasing radial profile of aspect ratio (see figure 9 in Han et al. (2025)). We find that an aspect ratio with a radial profile of $\mathcal{H}\propto R^{-7/2}$ – as predicted by Equation (51) – is consistent with the fitting results of Han et al. (2025). However, due to the relatively large uncertainties reported in that study, we refrain from further discussion of this system, instead we simply note that it is a promising target for future modeling and planet-disk interaction studies once higher-resolution data become available. [^16] HD 15115 is one of the systems that will be targeted by the first ALMA large program dedicated to debris disks, “The ALMA survey to Resolve exoKuiper belt Substructures” (ARKS; PI: S. Marino, ID: 2022.1.00338.L). Conversely, for the disks with $\mathcal{H}(R)\approx{\rm cte}$ according to Han et al. (2025), one may derive constraints on the maximum $M_{d}/m_{p}$ as a function of $a_{p}$ using Equation (26), i.e., for the system to be in the planet-dominated regime (Section 4.1).
@@ 6.4 Limitations and Future Work
The above points certainly warrant further investigations. However, the major limitation of this work is the omission of the full effects of the disk gravity. Namely, for the sake of analytical progress, only the axisymmetric component of the disk’s back-reaction was considered, while both its non-axisymmetric counterpart and the self-gravity were neglected (Sections 3 and 4). We shall address these limitations in an upcoming work. For now, we note that the framework outlined in Section 3 is well-suited for studying the effects of the disk’s full back-reacting potential. We expect the disk’s non-axisymmetric torque to mediate the exchange of angular momentum with the planet. Indeed, as we show in Appendix B, provided the system features a secular resonance, the interactions may lead to the exponential damping of the planetary inclination due to what is known as resonant friction or secular resonant damping (Tremaine, 1998; Ward & Hahn, 2003; Hahn, 2007). This happens such that $I_{p}(t)\propto\mathrm{exp}(-D_{\rm inc}t/2)$; see Equations (65)–(67). In such cases, the forced planetesimal inclinations will no longer be time-independent (contrary to this work), and will also decay over time. Indeed, if, at best, $D_{\rm inc}t\gg 1$ (Equation B), planetesimal inclinations will undergo damped sinusoidal oscillations that asymptotically converge to their free values, i.e., $I_{d}(t)\rightarrow I_{m}/2=\langle I_{d}(t)\rangle_{t}$ given by Equation (20). [^17] A similar behavior, though in the context of eccentricity dynamics, has been studied by Sefilian et al. (2023); see also Poblete et al. (2023). In such cases, the convergence $\mathcal{H}(R)\rightarrow\langle I_{d}(t)\rangle_{t}$ would happen faster, damping out the oscillatory behavior evident in Figure 8. Relatedly, the disk’s non-axisymmetric perturbations may also force the planetary inclination to oscillate over time, affecting the disk structure. Nevertheless, the dynamical regimes and solutions analyzed here remain generally valid.

### v2

@@ 5.2 Disks Less Massive than the Planet ($M_{d}/m_{p}\lesssim 1$)
Stage 1 $0\leq t\lesssim T_{\dot{\Omega}_{p}}]$: At early times, i.e., before the planet completes one full nodal precession cycle, the disk’s morphological evolution qualitatively resembles that of a massless disk. Indeed, as seen in Figures 9(i) and (ii), a spiral bending wave is launched at $a_{\rm in}$, which propagates outward as a leading pattern while gradually wrapping around the star. This manifests as a double-armed spiral in the top-down images and as a warp in the edge-on views, consistent with “Stage 1” evolution for $M_{d}=0$ disks (Section 5.1). However, this description is valid only within a certain radial range and is subject to minor modifications, as outlined below.
@@
Stage 2 $t\sim T_{\dot{\Omega}_{p}}]$: By the time the planet has completed one full precession cycle, the warp extends to $a_{\rm res}$, and the disk’s morphology exhibits significant deviations from that of a massless disk; see Figure 9(iii). First, the inner tilted region of the disk no longer transitions smoothly into the outer parts but instead does so rather abruptly. This creates the appearance of a “broken” disk, in which the region interior to the resonance appears separated from the outer parts. This can be understood by noting that planetesimal inclinations are maximized when nodally aligned with the planet in the inner region, but anti-aligned in the outer region (Figure 5(b)). This disconnect is most evident around the resonance, where inclinations are higher, especially compared to the inclinations toward the disk’s outer edge (recall that $I_{d}(a_{d})\propto a_{d}^{-7/2}$ at $a_{d}\gtrsim a_{\rm res}$; Equation 32).
@@
Stage 3 $t\gtrsim T_{\dot{\Omega}_{p}}]$: Further into the evolution, the disk settles into and maintains a quasi-equilibrium state, remaining largely unaffected by the continued growth of inclinations at the resonance; see panels (iv)–(vi) in Figure 9. Indeed, the disk structure – including the curly “X”-pattern – remains invariant, coprecessing with the planet’s ascending node as a rigid body. As this occurs, the broken, warp-like feature stalls near $a_{d}\approx a_{\rm res}$, persisting as a long-lived feature.
@@ 5.3 Disks More Massive than the Planet ($M_{d}/m_{p}\gtrsim 1$)
Stage 1 $0\leq t\lesssim T_{\dot{\Omega}_{p}}]$: At early times, the disk departs from its initial razor-thin configuration by developing a warp-like feature near its inner edge at $a_{\rm in}=30$ au (Figure 10(i)). This feature, however, is not directly analogous to the warp that develops in massless disks (Section 5.1). The key difference is that the location of the warp does not propagate outward over time. Instead, the maximum vertical displacement from the initial midplane gradually increases while remaining at an in-plane radius of $\approx a_{\rm in}$, such that $|Z_{\rm max}|\sim\sin[I_{d}(a_{\rm in})]$ (Figures 10(i)–(iii)). This effect is amplified as $I_{d}(a_{\rm in})\rightarrow 90^{\circ}$ in this specific setup (Figure 5(c)), although similar behavior would occur even if $I_{d}(a_{\rm in})$ were smaller. [^12] For instance, for a system with a larger $M_{d}/m_{p}$ and/or a smaller $a_{p}/a_{\rm in}$, such that $A_{d,p}/A_{p}$ at $a_{d}\approx a_{\rm in}$ is relatively larger than unity, rather than comparable to as in Model C; see the green curve in Figure 2. Additionally, this structure precesses as a whole while maintaining its symmetry around the planetary orbit.
@@
Finally, at $a_{d}\gtrsim a_{\rm in}$, the disk remains mostly undisturbed, preserving its flat structure. This is because, although planetesimals may have already attained their maximum inclinations, these decrease steeply with distance and remain small, $I_{m}(a_{d})\propto a_{d}^{-7/2}\ll 1^{\circ}$, with $\Delta\Omega(a_{d})\approx$ $\rm{const}$; see, e.g., panel (ii) in Figures 5(c) and 10.
Stage 2 $t\gtrsim T_{\dot{\Omega}_{p}}]$: Once the planet has completed more than one precession cycle, the warp-like feature begins to propagate outward, akin to the scenario with $M_{d}=0$ (Figures 10(iv)–(vi)). This occurs as the planetesimal inclination profile $I_{d}(a_{d})$ develops oscillations along $a_{d}$, and the ascending nodes $\Delta\Omega(a_{d})$ become phase-mixed in the inner disk regions; see also Figure 5(c). This description also explains the appearance of spiral patterns in the disk when viewed top-down, particularly near $a_{\rm in}$, where inclinations are larger.
@@ 5.4.4 Variation of the planetary semimajor axis, $a_{p}$
$\phi_{1}^{c}(a_{p}/a_{\rm in})\frac{M_{d}}{m_{p}}\sqrt{\frac{a_{\rm in}}{a_{p}}}={\rm const}.$ (45)
@@ 6 Discussion
To address the questions raised in Section 1.3 regarding the extent to which planetary inferences are compromised by the disk mass, we first discuss the implications of our findings for interpreting warped structures (§6.1), followed by vertical scale heights and density profiles (§6.2). Finally, we discuss the implications of our study for constraining the total masses of debris disks (§6.3), before reviewing the key assumptions and limitations of our model (§6.4).
We preface this section by emphasizing that our goal is not to provide definitive predictions for planet and disk parameters in any specific observed system. Rather, our aim is to demonstrate that even at the level of simplicity considered here – i.e., accounting only for the axisymmetric back-reaction of the disk on the planet, while neglecting its non-axisymmetric counterpart and the disk’s self-gravity – the resulting dynamics already alters the interpretation of debris disk observations, including planetary inferences, compared to models that assume massless disks. Thus, the quantitative predictions presented below should be viewed as illustrative rather than definitive. They serve as both a guide and motivation for future, more comprehensive studies that account for the full gravitational potential of the disk, whether via direct N-body integrators or specialized orbit-averaged codes (e.g., Hahn, 2003; Touma et al., 2009; Sefilian et al., 2023).
@@ 6.1.2 Application to HD 110058
We first consider HD 110058 (HIP 61782), which, in some ways, is reminiscent of $\beta$ Pictoris. This system, located at a distance of $130\pm 1$ pc (Gaia Collaboration et al., 2018), is part of the Lower Centaurus-Crux (LCC) association, with an estimated age of $15\pm 3$ Myr (Pecaut & Mamajek, 2016). It features a late A-type star with a dynamical mass of $\approx 1.84\pm 0.16M_{\odot}$ (Hales et al., 2022), and a near edge-on debris disk but no known planets (Meshkat et al., 2015; Brandt, 2021; Stasevic et al., 2023). The disk has been observed by multiple instruments (Kasper et al., 2015; Esposito et al., 2020; Stasevic et al., 2023; Ren et al., 2023), including ALMA (Lieman-Sifry et al., 2016; Hales et al., 2022), according to which the mm-sized dust grains populate the region between $\sim 18$ au and $70$ au. Interestingly, the disk appears to be vertically thick ($\mathcal{H}\approx 0.13$$0.28$; Hales et al., 2022) and warped, though the exact warp location is not well constrained, with values reported between $40$ and $60$ au (Kasper et al., 2015; Hales et al., 2022; Stasevic et al., 2023).
@@
Figure 11: The mass of a hypothetical planet $m_{p}$ as a function of its semimajor axis $a_{p}$ capable of producing a warp in the HD 110058 debris disk at $a_{\rm warp}=50$ au. The solution with a massless disk is depicted by the blue dashed curve ($M_{d}=0$). For the massive disks, two families of solutions are presented (assuming $p=3/2$; Eq. (1) and Section 6.1.1): the black and gray contours correspond to “Family 1” (i.e., $a_{\rm warp}=a_{\rm res}$) and “Family 2” (i.e., $a_{\rm warp}<a_{\rm res}$), respectively, with different contours representing different values of $M_{d}$ (in $M_{\earth}$; as indicated). Color-shaded regions are excluded: (i) the gray region represents where the disk would be too massive; (ii) the green region corresponds to inclination excitation timescales at the resonance that are longer than the system age; (iii) the yellow region indicates where the planet would have carved an inner disk edge larger than the one observed (Eq. 3); and (iv) the olive region represents where $m_{p}\geq M_{c}/10$. The remaining white (unshaded) area represents the allowed parameter space where all of the above conditions are satisfied. Note that the red region is where the presence of companions is ruled out with $68\%$ confidence (based on direct imaging and Gaia astrometry; Stasevic et al. (2023)). See the text (§6.1.2) for details.
@@
The results of Figure 11 are promising, in the sense that (planetary) companions responsible for the warp in HD 110058 could be more massive than currently thought (especially at $\sim 10$ au separations); therefore, potentially detectable in future observations. Currently, direct imaging and proper motion anomaly have ruled out giant planets outside the HD 110058 disk (namely, with $m_{p}\gtrsim 8M_{J}$ at $\gtrsim 50$ au; Brandt (2021); Wahhaj et al. (2013); Meshkat et al. (2015); Kervella et al. (2022)), and there is a likelihood of $68\%$ of detecting any companion interior to the disk with a mass smaller than $\sim 50M_{J}$ (Stasevic et al., 2023, see the red-shaded region in Figure 11).
@@ 6.1.3 The case of $\beta$ Pictoris
Figure 13: Vertical characteristics of the debris disks in each of models A, B, and C (Table 1). For each model, results correspond to the final snapshot time shown in Figure 5. Panel (A): Mass-weighted distributions of planetesimal inclinations, computed using Equation (18) and $p=3/2$, shown as histograms with $100$ equal-sized bins. For reference, the black vertical line represents twice the planetary inclination $2I_{p}(0)$. Note that in Model C, nearly $78\%$ of planetesimals have inclinations of $1^{\circ}$ or less. Panel (B): Normalized vertical density profiles as a function of distance from the disk midplane $Z_{\rm mid}$. Results are averaged over a radial range of $\pm 2.5$ au centered on the reference projected radius $R_{0}\approx 49.1$ au (equal to $\sqrt{a_{\rm in}a_{\rm res}}$ in Model B), using a resolution of $dZ_{\rm mid}=0.1$ au. The inset presents a zoomed view around the midplane. Regardless of disk mass, the non-Rayleigh nature of the inclination distributions and the non-Gaussian shape of the density profiles are evident. For $M_{d}/m_{p}=0.1$, the inclination distribution is bimodal (apart from the peak at $I_{d}=90^{\circ}$), and the density profile resembles a sum of two Gaussians with different widths, indicative of dynamically cold and hot populations. See the text (Section 6.2.1) for details.
@@ 6.2.2 Aspect ratios: radial dependence and implications
Second, the radial profile of a disk’s aspect ratio – whether constant or decreasing with $R$ – can provide insights into the effects of disk gravity. In principle, this could constrain the parameters of a potential planet as a function of $M_{d}$ by making use of Equations (49)–(51), or conversely, offer a dynamical estimate of the disk mass if a planet is already known. However, by and large for simplicity, most model fits to observations typically assume a Gaussian vertical profile with a distance-independent aspect ratio, $\mathcal{H}(R)={\rm const}$ (e.g., Kennedy et al., 2018; Daley et al., 2019; Vizgan et al., 2022; Terrill et al., 2023; Matrà et al., 2025). As a result, this approach cannot currently be pursued, and future modelling efforts should aim to account and fit for the radial dependence of vertical scale heights.
We note that a recent study by Han et al. (2025) explored this avenue using a non-parametric fitting procedure. They found that 8 out of 9 considered disks are consistent with a distance-independent aspect ratio or an $\mathcal{H}(R)$ which increases with radius in their millimeter dust continuum. Interestingly, the debris disk around HD 15115 (MacGregor et al., 2019, not known to have companions) stands out due to its decreasing radial profile of aspect ratio (see figure 9 in Han et al. (2025)). We find that an aspect ratio with a radial profile of $\mathcal{H}\propto R^{-7/2}$ – as predicted by Equation (51) – is consistent with the fitting results of Han et al. (2025). However, due to the relatively large uncertainties reported in that study, we refrain from further discussion of this system, instead we simply note that it is a promising target for future modeling and planet-disk interaction studies once higher-resolution data become available. [^16] HD 15115 is one of the systems that will be targeted by the first ALMA large program dedicated to debris disks, “The ALMA survey to Resolve exoKuiper belt Substructures” (ARKS; PI: S. Marino, ID: 2022.1.00338.L). Conversely, for the disks with $\mathcal{H}(R)\approx{\rm const}$ according to Han et al. (2025), one may derive constraints on the maximum $M_{d}/m_{p}$ as a function of $a_{p}$ using Equation (26), i.e., for the system to be in the planet-dominated regime (Section 4.1).
@@ 6.4 Limitations and Future Work
The above points certainly warrant further investigations. However, the major limitation of this work is the omission of the full effects of the disk gravity. Namely, for the sake of analytical progress, only the axisymmetric component of the disk’s back-reaction was considered, while both its non-axisymmetric counterpart and the self-gravity were neglected (Sections 3 and 4). We shall address these limitations in an upcoming work. For now, we note that the framework outlined in Section 3 is well-suited for studying the effects of the disk’s full back-reacting potential (but not the disk self-gravity).
Our preliminary numerical calculations using the general form of Equation (15) show that the primary effect of the disk’s non-axisymmetric torque on the planet is to mediate the exchange of angular momentum with the planet. Indeed, as we show in Appendix B, provided the system features a secular resonance, the interactions may lead to the exponential damping of the planetary inclination due to what is known as resonant friction or secular resonant damping (Tremaine, 1998; Ward & Hahn, 2003; Hahn, 2007). This happens such that $I_{p}(t)\propto\mathrm{exp}(-D_{\rm inc}t/2)$; see Equations (65)–(67). In such cases, the forced planetesimal inclinations will no longer be time-independent (contrary to this work), and will also decay over time. Indeed, if, at best, $D_{\rm inc}t\gg 1$ (Equation B), planetesimal inclinations will undergo damped sinusoidal oscillations that asymptotically converge to their free values, i.e., $I_{d}(t)\rightarrow I_{m}/2=\langle I_{d}(t)\rangle_{t}$ given by Equation (20). As a result, the distribution of $\Delta\Omega(a_{d})$ may eventually span the full $\pi,\pi]$ range in both planet- and disk-dominated regimes. [^17] A similar behavior, though in the context of eccentricity dynamics, has been studied by Sefilian et al. (2023). In such cases, for instance, the convergence of $\mathcal{H}(R)\rightarrow\langle I_{d}(t)\rangle_{t}$ would happen faster, damping out the oscillatory behavior evident in Figure 8. Relatedly, the disk’s non-axisymmetric perturbations may also force the planetary nodal precession rate to deviate from that given by Equation (27); our preliminary calculations suggest that the deviation could be by a factor of a few, thus potentially affecting inferences of, e.g., disk masses by a similar factor. Nevertheless, we find that the dynamical regimes and solutions analyzed here remain generally valid, giving us confidence that the simplified analytical framework – together with Appendix B – adequately capture the basic effects of a fully back-reacting disk.
This said, we note that the influence of disk self-gravity remains an open question requiring further investigation. Silsbee & Rafikov (2015) investigated the in-plane dynamics of planetesimals around eccentric stellar binaries by modeling the massive protoplanetary disk as a rigid, precessing slab and demonstrated the existence of up to four distinct dynamical regimes. These regimes arise depending on whether the disk or the inner companion dominates the precession rates and eccentricity excitation of planetesimals, with all combinations theoretically feasible (see also Sefilian & Touma, 2019). Analogous dynamical complexity for inclinations may also manifest in gravitating debris disk–planet systems. We hope that our analytical results will provide a foundation and starting point for future investigations of these phenomena, whether through direct $N$-body simulations or secular “$N$-ring" codes (e.g., Hahn, 2003; Touma et al., 2009; Sefilian et al., 2023).